A shaft fitted with a flywheel rotates at 250 rpm and drives a machine the resisting torque of


which varies in a cyclic manner over a period of three revolutions. The torque rises from 675 Nm to


2700 Nm in a uniform manner during 0.5 revolution and remains constant for 1 revolution, the cycle


being then repeated. If the driving torque applied to the shaft is constant and the flywheel has a


mass of 450 kg and a radius of gyration of 0.6 m, find the power necessary to drive the machine and


the percentage fluctuation of speed.

1
Expert's answer
2021-11-04T00:41:43-0400

Power of the motor

P=Mean torque*Angular speed

"=T_m*\frac{2πN}{60}"

"=1875*\frac{2π*250}{60}"

"=49.0625W"

Maximum fluctuation energy is given by "∆E=\frac{1}{2}*\frac{π}{2}*1125+(2π*1125)+\frac{1}{2}*\frac{π}{2}*1125"

"=1125(\frac{π}{4}+\frac{π}{4}+2π)=1125(\frac{π}{2}+2π)"

"=\frac{5π}{2}*1125=8831.25"

"∆E=mk^2\omega^2C_s"

"8831.25=450(0.6)^2*(26.166)^2*C_s"

"8831.25=110914.8481C_s"

"C_s=0.079622"

Hence percentage fluctuation of speed =7.9622%